# Disentangling $\gamma-\beta$: the 4th-order velocity moments based on spherical Jeans analysis

### v1

@@ 5 Conclusions
We construct a dynamical model for dwarf spheroidal and ultra-faint dwarf galaxies under the frame of 2nd-order and 4th-order spherical Jeans equations. Assuming that the velocity anisotropy $\beta$ is constant with the galactocentric distance and negligible stellar mass compared to dark mass, the model is examined using two sets of mock galaxies, one cuspy and the other cored, to investigate its behavior and ability to break mass-anisotropy degeneracy. To show the effects of some observational limitations, we have performed fittings with different values of a number of sample stars, l.o.s. velocity error measurements, and spatial distribution of the sample. We also conduct fittings where only 2nd-order moments are taken into account.
The results of our fitting show that the incorporation of 4th-order moments disentangles $\beta$ from $\gamma$. It partially reduces the degeneracy between the two parameters compared to the situation when only 2nd-order moments are taken into account. The heavy reliance of 4th-order moments on the tails of the velocity distribution, where the star count is low, makes it prone to velocity error measurements. We find that in this model, $\beta$ is consistently recovered within $1\sigma$ uncertainty as long as the ratio $\sigma_{\rm los,global}/\Delta v_{\rm los}\gtrsim 4$ is satisfied, even when the sample size is as small as 500 stars. While the ratio $\sigma_{\rm los,global}/\Delta v_{\rm los}$ holds the most control on the systematic bias, the number of stars affects the most on the tightness of posterior distribution. However, even though this model succeeds in reproducing the input value of $\beta$, the condition for $\gamma$ is not as simple. The successful recovery of $\beta$ is mainly aided by $\kappa_{\rm los}(R)$’s sensitivity to the change in $\beta$. The milder response of $\kappa_{\rm los}(R)$ to changes in $\gamma$ results in a less strict constraint on $\gamma$. We find that the incorrect value of $\gamma$ can be ruled out more robustly as $\kappa_{\rm los}(R)$ deviates more from 3. This happens because of the existence of degeneracy between $\gamma-a_{\rm dm}-\rho_{0}$ and the fact that if $\kappa_{\rm los}$ is far from 3, any change of these parameters changes $\kappa_{\rm los}$ more dramatically compared to when $\kappa_{\rm los}\approx 3$.
@@
The challenging situation when $\kappa_{\rm los}\approx 3$ indicates the importance of relaxing the assumption of sphericity to step forward and adopt a more realistic condition. In an axisymmetric model, for example, different l.o.s. velocity dispersion profiles are imprinted along the projected major and minor axes with also distinguishable responses between a core and a cusp (Hayashi & Chiba, 2012). Although the behavior of the l.o.s. kurtosis in the axisymmetric model remains unexplored, incorporating these signatures is expected to enable more stringent estimations to uncover the actual density profile in dwarf galaxies. Moreover, the support for the non-spherical model is reinforced by both observations and simulations. It is noted that the light distribution in dSphs and UFDs is not spherically symmetry (Muñoz et al., 2018; Battaglia et al., 2022). From the theoretical side, results from CDM simulations show that dark halos are generally non-spherical (Allgood et al., 2006; Orkney et al., 2023).
@@
In this work, we only consider conditions where the membership of a star can be determined perfectly. However, working with observational data has to face the challenge of contamination from foreground or unbound stars (Battaglia et al., 2013). While a detailed study of contamination effects is beyond the scope of this paper, the use of kurtosis is more prone to these contaminations compared to the model using only the 2nd-order moments, especially if those stars fall in the tails of LOSVDs (Merrifield & Kent, 1990). The role of binaries also needs to be properly taken into account. Although it is suggested that the contribution of binaries in inflating the velocity dispersion profile in dSphs is insignificant and to be carefully treated if we work with UFDs (Minor, 2013; Spencer et al., 2017; Arroyo-Polonio et al., 2023; Wang et al., 2023), a quantitative approach of unresolved binaries influence on $\kappa_{\rm los}(R)$ is to be studied in future work.
@@
We are grateful to Laszlo Dobos, Carrie Filion, Evan Kirby, Alex Szalay, Rosie Wyse, and the Subaru/PFS GA working group for their useful discussion on the derivation and analysis of the LOSVD of dwarf satellites. This work was supported in part by the MEXT Grant-in-Aid for Scientific Research (No. JP20H01895, JP21K13909, and JP23H04009 for KH, No. JP18H05437, JP21H05448, and JP24K00669 for MC.).
@@ A.2 Laplacian Kernel
$f_{s}=\frac{b}{4a}\exp\left(\frac{t^{2}-2abw}{2a^{2}}\right)\operatorname{erfc}{\left(\frac{t^{2}-abw}{\sqrt{2}ta}\right)}+\\ \frac{b}{4a}\exp\left(\frac{t^{2}+2abw}{2a^{2}}\right)\operatorname{erfc}{\left(\frac{t^{2}+abw}{\sqrt{2}ta}\right)},$ (A9)
@@
with the same relation as Equation A3, A4, A5 for $t$, $b$, $k$, respectively. The value for some appearing constants are $a_{0}=2.25,k_{0}=1/\sqrt{2},k_{\infty,0}=1.08$. This kernel can produce distributions with kurtosis lying in the range of $3.0<\kappa\leqslant 6.0$.
Appendix B All free parameters
[…]

### v2

@@ 5 Conclusions
We develop a dynamical model for dwarf spheroidal and ultrafaint dwarf galaxies using the 2nd-order and 4th-order spherical Jeans equations. In this model, we assume a constant velocity anisotropy parameter, $\beta$, with respect to galactocentric distance, and we consider the stellar mass to be negligible compared to the dark matter mass. The model is tested on four sets of mock galaxies to assess its performance in breaking the mass-anisotropy degeneracy and distinguishing between cored and cuspy density profiles. To examine the impact of observational limitations, we perform fits using different sample sizes, line-of-sight (l.o.s.) velocity measurement errors, and spatial distributions of stars. Additionally, we conduct fits that consider only the 2nd-order moments.
The results of our fitting demonstrate that incorporating 4th-order moments effectively disentangles $\beta$ from $\gamma$, partially breaking the degeneracy between these two parameters compared to the scenario where only 2nd-order moments are taken into account. We find that, in this model, a ratio of $\sigma_{\rm los,global}/\delta v_{\rm los}\gtrsim 4$ is necessary to mitigate most of the systematic biases associated with velocity errors. While this ratio plays the dominant role in controlling systematic bias, the number of stars has the greatest influence on the precision of the posterior distributions. Additionally, the results also suggest that the incorrect value of $\gamma$ can be ruled out more robustly as $\kappa_{\rm los}(R)$ deviates further from 3. Specifically, due to the nature of a cuspy density profile, which restricts $\kappa_{\rm los}(R)$ to a relatively narrow range compared to a cored profile, ruling out a cuspier profile than the true underlying one is less challenging than ruling out shallower density profiles.
@@
While this model assumes spherical symmetry, real systems are often more complex and, thus, require a broader framework. Both observations and simulations indicate support for non-spherical models. Observationally, the light distribution in dSphs and UFDs is not spherically symmetric (Muñoz et al., 2018; Battaglia et al., 2022). On the theoretical side, CDM simulations reveal that dark matter halos are generally non-spherical (Allgood et al., 2006; Orkney et al., 2023). In particular, studies of axisymmetric models have shown that distinct line-of-sight velocity dispersion profiles are projected along the major and minor axes, with notable differences between core and cusp profiles (Hayashi & Chiba, 2012). Although the behavior of line-of-sight kurtosis in axisymmetric models remains unexplored, incorporating these features is expected to allow a more precise estimation of the true density profile in dwarf galaxies.
@@
Despite these considerations regarding galaxy structure, our analysis in this study assumes idealized conditions, where the membership of each star can be determined with perfect accuracy. However, observational data often suffer from contamination by foreground or unbound stars (Battaglia et al., 2013). Although a detailed analysis of contamination effects is beyond the scope of this paper, it is worth noting that kurtosis-based models are generally more sensitive to such contamination than models using only 2nd-order moments, especially if the contaminations lie in the tails of the LOSVD (Merrifield & Kent, 1990). Additionally, the influence of binary stars needs to be carefully considered. While the contribution of binaries to inflating the velocity dispersion profile in dSphs is demonstrated to be minimal (Minor, 2013; Spencer et al., 2017; Arroyo-Polonio et al., 2023; Wang et al., 2023), their effect could be more significant in UFDs and warrants careful treatment. A quantitative analysis of how unresolved binaries impact $\kappa_{\rm los}(R)$ will be an important focus of future work.
@@
We thank the anonymous reviewer for constructive comments that helps revise the paper. We are also grateful to Laszlo Dobos, Carrie Filion, Evan Kirby, Alex Szalay, Rosie Wyse, and the Subaru/PFS GA working group for their useful discussion on the derivation and analysis of the LOSVD of dwarf satellites. This work was supported in part by the MEXT Grant-in-Aid for Scientific Research (No. JP20H01895, JP21K13909, and JP23H04009 for KH, No. JP18H05437, JP21H05448, and JP24K00669 for MC.).
@@ A.2 Laplacian Kernel
$f_{s}(w)=\frac{b}{4a}\exp\left(\frac{t^{2}-2abw}{2a^{2}}\right)\operatorname{erfc}{\left(\frac{t^{2}-abw}{\sqrt{2}ta}\right)}+\\ \frac{b}{4a}\exp\left(\frac{t^{2}+2abw}{2a^{2}}\right)\operatorname{erfc}{\left(\frac{t^{2}+abw}{\sqrt{2}ta}\right)},$ (A9)
@@
with the same relation as Equations (A3), (A4), (A5) for $t$, $b$, $k$, respectively. The value for some appearing constants are $a_{0}=2.25,k_{0}=1/\sqrt{2},k_{\infty,0}=1.08$. This kernel can produce distributions with kurtosis lying in the range of $3.0<\kappa\leqslant 6.0$.
Appendix B Dependency on the prior of $\gamma$
The solutions generated from Jeans modeling are not always guaranteed to be physically plausible. This limitation becomes particularly relevant when selecting the prior range for $\gamma$ because solutions with $\gamma<0$ are unlikely to represent realistic physical conditions and can slow the convergence of MCMC chains. However, simply excluding $\gamma<0$ from the prior range may unintentionally introduce boundary effects.
To examine the impact of the prior choice on $\gamma$, we compare results using different prior ranges. Figure 12 presents the posterior distributions of each parameter derived from fitting 5,000 randomly selected stars in the cored dSph mock galaxies. We apply two different ranges for the prior of $\gamma$: a broader range of $-2.0\leqslant\gamma\leqslant 2.0$ (shown in blue) and a narrower range of $0\leqslant\gamma\leqslant 2.0$ (shown in orange). The red lines indicate the true values. The estimated median values and 95% confidence intervals for each case are also indicated above the corresponding column, as well as by three vertical dashed lines.
Except for the hard cut at $\gamma=0$, Figure 12 shows that the posterior distribution of $\gamma$ in the narrow-prior case closely resembles that of the broad-prior case. However, the other parameters must compensate for this restriction in $\gamma$ to maintain the fit to the data. These compensations propagate through the correlations between $\gamma$ and other parameters, leading to shifts, broadening, or artificial narrowing in their posterior distributions. This effect is, for example, pronounced for $\rho_{0}$, where $\gamma-\rho_{0}$ correlation causes the exclusion of the true value. To minimize these prior-induced effects, we apply the broader prior $-2.0\leqslant\gamma\leqslant 2.0$ to all our fittings.
Figure 12: Posterior distributions on $\gamma$ obtained from the fittings for the cored dSph mock galaxy by applying different ranges of prior. The blue posteriors correspond to a wide range of prior $-2.0\leqslant\gamma\leqslant 2.0$, and the orange posteriors are obtained by assigning a narrower prior range of $0\leqslant\gamma\leqslant 2.0$ that excludes unphysical values. The red solid lines denote the true value, while the three vertical dashed lines indicate the median and 95% confidence intervals.
Appendix C Posterior distribution for all parameters
Figure 13 and 14 presents the posterior distributions of all six free parameters in our analysis: $\beta$, $\alpha$, $\beta_{\rm trans}$, $\gamma$, $\rho_{0}$, and $b$, displayed in each panel from left to right. The left (right) column corresponds to the cuspy (cored) dSph mock galaxy. The number of stars and velocity errors are indicated on the left side for each row. All panels are obtained from a model that incorporates 4th-order moments. The red lines represent the true input values of each parameter, and the vertical dashed lines denote the median and 68% confidence level. Correlations between $b$, $\gamma$, and $\rho_{0}$ appear in the two-dimensional distributions, as highlighted by the diagonally elongated contours.
Figure 13: Posterior distributions of all free parameters from the fittings of the cuspy dSph mock galaxy (left column) and the cored dSph mock galaxy (right column). The number of stars and velocity error used in each case are indicated on the leftmost side of each row. All results are obtained using the model incorporating 4th-order moments. Red solid lines indicate the true parameter values, while vertical dashed lines in the 1-dimensional posterior distributions represent the median and 68% confidence intervals (note that these differ from the 95% confidence intervals shown in Figure 5). Estimated median values and 68% confidence intervals for each parameter are also indicated above the corresponding columns. Contour lines in the 2-dimensional distributions represent the 68% and 95% confidence levels.
Figure 14: Continuation of Figure 13.
Appendix D Standard deviation of the medians among realizations
Table 6 summarizes the standard deviations of $\gamma$ and $\beta$ across multiple realizations for each fitting scenario applied to the dSph mock galaxies. These standard deviations provide a quantitative measure of the variability in the recovered parameters between different realizations.
Table 6: Standard deviation of the medians across realizations from the results of dSph mock galaxies fitting.
Parameter Density profile $N=5,000$ $N=5,000$ $N=500$
$\delta v_{\rm los}=0.01$ km s-1 $\delta v_{\rm los}=2.00$ km s-1 $\delta v_{\rm los}=2.00$ km s-1
2nd 4th 2nd 4th 2nd 4th
$\gamma$ cuspy $0.19$ $0.09$ $0.13$ $0.24$ $0.42$ $0.31$
cored $0.25$ $0.18$ $0.27$ $0.71$ $0.54$ $0.50$
$\beta$ cuspy $0.08$ $0.07$ $0.09$ $0.16$ $0.30$ $0.19$
cored $0.14$ $0.07$ $0.16$ $0.13$ $0.63$ $0.26$